Allison:

An investor has an account with stock from two different companies. Last year, her stock in Company A was worth $2100 and her stock in Company B was worth $5120. The stock in Company A has decreased 9% since last year and the stock in Company B has decreased 5%. What was the total percentage decrease in the investor's stock account? Round your answer to the nearest tenth (if necessary).

AZ:

Ugh it's going to be a pain since I can't type replies with enters

AZ:

so the investor has two accounts, Company A has 2100 and decreased by 0.09 so essentially, what is the value of it now?

AZ:

It'll be \( 2100 \times (1 - 0.09)= ?\)

AZ:

and then company B stocks were worth 5120 and they decreased by 5% so \( 5120 \times (1-0.05) = ?\)

AZ:

That's how you know how much it decreased too. Basically, you're decreasing the first one by 9% so 0.09 * 2100 and then you have to subtract that number from 2100 and so I just re-wrote that as 2100 * (1-0.09) so that way you can save time and steps since you get the same answer both ways

AZ:

and then after that, add those two numbers you calculated and then you have to use this formula

AZ:

Your final number is the two numbers you added together after calculating the new value of the stocks. Your initial value is going to be the original numbers added together so 2100 + 5120

AZ:

\(\text{% decrease} = \dfrac{\text{final - initial}}{\text{initial}}\times 100\)

Allison:

Oh my God what is this

AZ:

@az wrote:
It'll be \( 2100 \times (1 - 0.09)= ?\)
One step at a time, can you first tell me what this is